OpenStudy (anonymous):

how do you solve this equation \sqrt{9-b}=\sqrt{1-9b}

OpenStudy (anonymous):

square both sides

OpenStudy (anonymous):

i tried dont get it

OpenStudy (anonymous):

What do you get?

OpenStudy (anonymous):

80

OpenStudy (anonymous):

9-b = 1-9b is what you should get after you square it

OpenStudy (anonymous):

yeah i dont get it sorry

OpenStudy (anonymous):

what is sqrt(5) * sqrt(5) ?

OpenStudy (anonymous):

25

OpenStudy (anonymous):

\[\sqrt{x^2} = (\sqrt{x})^2 = x\]

OpenStudy (anonymous):

whats that